X company has a demand of 75 Lakhs per year. Setup cost and holding cost are 29 and 35 per unit. Number of pieces in a order includes 2.3 Lakhs per order. Printing is done 300 days and 2 days to deliver this magazine.What can be the EOQ of this? With above data find the Setup Cost and Holding cost for a year. What is the Reorder Point. What is the annual demand to production ratio in this case. What does it indicate for
1
Expert's answer
2021-07-12T17:41:01-0400


Given that

Annual demand = 75 lakhs per year

Setup cost = $29 per unit

Holding cost = $35 per unit

Number of pieces in one order = 2.3 lakhs

Working days = 300 days

Lead time = 2 days

The formula to calculate EOQ is


EOD= "\sqrt{2SD \over H}"


Here, S = order cost

D = annual demand

H = holding cost

Substituting the values,


EOD= "\sqrt{2 \times 29 \times7500000 \over 35}"


EOD= "\sqrt{12428571.42}"


EOD= 3525.41